Customers of a phone company can choose between two service plans for long distance calls. The first plan has a $35 one-time activation fee and charges 7 cents a minute. The second plan has no activation fee and charges 12 cents a minute. After how many minutes of long distance calls will the costs of the two plans be equal?

Answer by mananth(16949)

Customers of a phone company can choose between two service plans for long distance calls. The first plan has a $35 one-time activation fee and charges 7 cents a minute. The second plan has no activation fee and charges 12 cents a minute. After how many minutes of long distance calls will the costs of the two plans be equal?

Let the number of minutes after which the they become equal be x
Convert call charges from cents to dollars
first plan will cost 0.07x + 35
Second plan will cost 0.12x
0.07x +35 = 0.12x
0.07x-0.12x= -35
-0.05x=-35
x=-35/-0.05
x=700 minutes
after 700 minutes they will become equal